# ECG-Soup: Harnessing Multi-Layer Synergy for ECG Foundation Models

### v2

@@ 4.1 Theoretical analysis
where are random embeddings of masked patches.
@@ 5.3 Implementation Details
In particular, within the generative learning group, we implemented and evaluated the spatio-temporal Masked Vision Transformer (STMEM), which also served as the foundation for our proposed self-supervised model. In generative learning, the STMEM and our proposed methods used the ViT backbone, where the ECG signal was partitioned into non-overlapping segments (patches) with a patch size of 50 signal points each. During pretraining, 75% of the patches are masked randomly, and the model was trained to reconstruct the masked patches by minimizing the MSE loss function between the original and reconstructed patches. The implementation details are given in the Table III.
@@ 6.2 Correlation between patch embeddings:
The comparison between the two plots in Figure 3 reveals a significant difference in the representation organization. STMEM exhibits a gradual convergence process, with patch similarity steadily increasing across layers before dropping sharply in the final layers. IPASTMEM achieves a high similarity early on (layers 4–5) and then gradually declines. This suggests that IPASTMEM tends to “compress” information at the beginning of the network, then re-establishes representational diversity in later layers. Notably, these properties are consistently maintained across all ECG leads, confirming that the model’s representation learning mechanism does not depend on the individual characteristics of each signal channel but instead exhibits a generalization tendency across the entire input data space. Thus, pretrained ViT models in multi-lead ECG signal processing demonstrate high stability and adaptability. The results of the average cosine similarity across layers explain the phenomenon observed in the linear probing evaluation across layers: the middle layers often offer superior performance for downstream tasks. The reason is that the model reaches an optimal information synthesis state at this stage, in which the representations effectively encode the specific spatio-temporal relationship of the ECG signal.
@@ 6.3 Average attention entropy:
Analyzing the average attention entropy (AAE) across layers, as shown in Figure 5, provides additional evidence to explain the performance degradation on downstream classification tasks when using the representation from the last layer. Both models exhibit very high entropy in the first layer, reflecting the dispersion of attention and the high uncertainty about the input information. This is typical of the early stages of representation learning when the model has not yet identified the critical regions in the signal. From the second layer onward, the entropy drops rapidly. It stabilizes at a low level, indicating that the model begins to focus on meaningful information and the representation becomes more certain. In the final layers, the entropy increases slightly again, indicating that the model pays more attention to the fine-grained features in the ECG signal. However, the entropy level at the last layer is still significantly lower than at the first. When these observations are compared with the results from linear probing and cosine similarity, it can be concluded that the final layers are not a crude representation but rather a reconstruction stage designed to focus on high-resolution features. The mathematical definition of AAE is provided in Equation 8, 9.
@@ 7 Experiments and Results
We evaluated the ECG representations learned by the various methods using two popular strategies: linear probing and fine-tuning, applied in both in-distribution and OOD settings. Experimental analyses demonstrated that the proposed methods significantly improve most evaluation metrics, indicating superior performance to the baselines.
@@ 7.3 Ablation Study
The combined analysis from Tables VII, VIII, and IX (Appendix D) shows a clear difference in the performance gap between linear probing and full model fine-tuning across the method groups. SimCLR consistently exhibits a huge gap: relatively low performance with linear probing but a sharp increase with fine-tuning on both PTB-XL (in-distribution and OOD) and Chapman, reflecting that the representations learned from SimCLR are not informative enough and rely almost entirely on full parameter updates to achieve high performance. In contrast, CPC and STMEM exhibit only a small gap, indicating stable representation quality and better generalization ability. In particular, the proposed methods maintain a small gap between the two settings and, in some cases, slightly degrade performance, indicating that they have learned informative and stable representations robust enough to achieve high performance with linear probing alone. However, when applying fine-tuning, the proposed methods maintain superior performance over the entire baseline, while the observed characteristics in Tables IV, V, VI (linear probing) remain unchanged.
@@ A.2 Masked Vision Transformer
The ECG patches are randomly masked in the pre-text task. The unmasked patches are fed into the ViT encoder, while the masked patches are reconstructed through the decoder-based Transformers as illustrated in Figure 6. Encoder: In the pretraining phase, a random masking strategy is applied on the embedding sequence to reduce the inherent redundant information in the ECG signal and simultaneously avoid overfitting. We denote the unmasked embeddings as
@@ Appendix B The proof of lemma 1
If , then . 3 holds trivially. Consider and then normalizing , :
Applying the norm of both sides of 32 yields
@@ Appendix C Hidden layer analysis
Figure 7 illustrates the classification performance of the STMEM and IPASTMEM models on each representation layer (from layer 1 to layer 12) on four datasets: PTB-XL Conditions, PTB-XL Rhythms, Chapman Conditions, and Chapman Rhythms. The plots show the macro AUC, sample AUC, instance precision, sample accuracy, macro F1 score, and sample F1 score (h) values when training the classifier on the representations extracted across layers in the pretrained ViT model.
@@ Appendix D Ablation study
| Condition | Self-supervised | SimCLR | 84.970.28 92.410.51 (+7.44) | 94.630.31 96.860.18 (+2.23) | 29.371.47 35.341.43 (+5.97) | 97.450.02 97.840.06 (+0.39) | 12.430.62 24.311.23 (+11.88) | 61.880.84 70.280.65 (+8.40) |
| CPC | 89.970.3 91.240.36 (+1.27) | 96.040.09 96.520.09 (+0.48) | 32.600.39 34.610.22 (+2.01) | 97.660.01 97.770.02 (+0.11) | 17.120.69 19.620.69 (+2.50) | 66.230.2 68.260.26 (+2.03) | ||
| STMEM | 92.600.17 93.690.28 (+1.09) | 96.920.07 97.410.1 (+0.49) | 34.780.6 38.261.06 (+3.48) | 97.820.01 98.020.03 (+0.20) | 21.60.9 26.211.43 (+4.61) | 68.410.35 72.40.6 (+3.99) | ||
@@
| Rhythm | Self-supervised | SimCLR | 87.120.28 96.270.6 (+9.15) | 96.280.48 97.320.49 (+1.04) | 81.80.36 89.050.78 (+7.25) | 97.430.04 98.530.08 (+1.1) | 25.821.39 44.882.48 (+19.06) | 83.770.63 89.890.96 (+6.12) |
| CPC | 91.510.37 92.980.74 (+1.47) | 94.330.17 94.580.14 (+0.25) | 82.050.23 83.210.23 (+1.16) | 97.740.03 97.910.04 (+0.17) | 30.692.58 36.52.34 (+5.81) | 82.250.27 83.290.23 (+1.04) | ||
| STMEM | 97.250.21 97.040.26 (-0.21) | 95.330.12 95.510.17 (+0.18) | 85.550.38 86.470.3 (+0.92) | 98.240.05 98.360.04 (+0.12) | 44.12.73 47.872.58 (+3.77) | 85.720.28 86.570.35 (+0.85) | ||
@@
| Condition | Self-supervised | SimCLR | 77.370.35 83.650.44 (+6.28) | 89.440.62 97.70.47 (+8.26) | 36.350.72 54.472.34 (+18.12) | 98.170.02 98.790.08 (+0.62) | 12.520.95 22.591.62 (+10.07) | 57.810.95 79.381.09 (+21.57) |
| CPC | 81.960.48 82.690.34 (+0.73) | 95.190.26 96.130.31 (+0.94) | 47.280.63 49.850.61 (+2.57) | 98.510.02 98.610.03 (+0.1) | 14.870.68 16.560.61 (+1.69) | 70.40.47 73.370.5 (+2.97) | ||
| STMEM | 85.240.22 85.370.25 (+0.13) | 97.760.24 98.220.23 (+0.46) | 53.580.66 55.421.02 (+1.84) | 98.770.01 98.840.03 (+0.07) | 19.761.12 21.820.95 (+2.06) | 78.50.52 80.390.77 (+1.89) | ||
@@
| Rhythm | Self-supervised | SimCLR | 83.310.24 93.870.8 (+10.56) | 86.640.59 97.10.73 (+10.46) | 61.961.24 89.351.28 (+27.39) | 95.550.09 98.510.15 (+2.96) | 37.281.52 55.752.42 (+18.47) | 64.881.07 89.971.43 (+25.09) |
| CPC | 91.470.36 92.390.73 (+0.92) | 94.030.62 95.310.31 (+1.28) | 80.181.25 84.040.75 (+3.86) | 97.320.11 97.780.13 (+0.46) | 45.810.7 47.961.22 (+2.15) | 81.21.12 84.660.8 (+3.46) | ||
| STMEM | 94.420.07 94.40.13 (-0.02) | 97.60.16 97.90.24 (+0.3) | 90.440.47 91.351.02 (+0.91) | 98.670.06 98.750.14 (+0.08) | 57.990.88 59.041.76 (+1.05) | 91.160.29 91.860.87 (+0.7) | ||
@@
| Condition | Self-supervised | SimCLR | 84.650.19 92.380.42 (+7.73) | 94.690.22 96.960.19 (+2.27) | 28.891.13 34.751.51 (+5.86) | 97.470.01 97.850.04 (+0.38) | 13.960.69 24.512.13 (+10.55) | 62.340.31 70.430.73 (+8.09) |
| CPC | 89.70.44 90.640.27 (+0.94) | 95.910.15 96.350.07 (+0.44) | 31.930.25 33.80.31 (+1.87) | 97.620.01 97.740.01 (+0.12) | 15.880.94 19.330.39 (+3.45) | 65.30.39 67.750.22 (+2.45) | ||
| STMEM | 92.170.18 93.560.29 (+1.39) | 96.890.12 97.430.09 (+0.54) | 34.180.36 38.230.75 (+4.05) | 97.790.01 98.020.01 (+0.23) | 21.681.32 27.011.06 (+5.33) | 68.140.31 72.570.43 (+4.43) | ||
@@
| Rhythm | Self-supervised | SimCLR | 84.980.71 96.440.45 (+11.46) | 96.170.36 97.410.43 (+1.24) | 81.380.6 89.090.73 (+7.71) | 97.360.05 98.540.09 (+1.18) | 24.420.87 44.282.87 (+19.86) | 83.320.43 90.060.82 (+6.74) |
| CPC | 90.420.81 92.111.37 (+1.69) | 94.250.11 94.40.07 (+0.15) | 81.280.4 82.220.24 (+0.94) | 97.590.05 97.730.05 (+0.14) | 32.213.04 33.892.37 (+1.68) | 81.540.32 82.370.16 (+0.83) | ||
| STMEM | 96.790.32 96.830.28 (+0.04) | 95.470.11 95.510.2 (+0.04) | 85.780.16 86.490.35 (+0.71) | 98.240.04 98.380.03 (+0.14) | 46.334.18 49.443.26 (+3.11) | 85.980.23 86.690.35 (+0.71) | ||

### v3

@@ 4.1 Theoretical analysis
where are learnable embeddings of masked patches.
@@ 5.3 Implementation Details
In particular, within the generative learning group, we implemented and evaluated the STMEM, which also served as the foundation for our proposed self-supervised model. In generative learning, the STMEM and our proposed methods used the ViT backbone, where the ECG signal was partitioned into non-overlapping segments (patches) with a patch size of 50 signal points each. During pretraining, 75% of the patches are masked randomly, and the model was trained to reconstruct the masked patches by minimizing the MSE loss function between the original and reconstructed patches. The implementation details are given in the Table III.
@@ 6.2 Correlation between patch embeddings:
The comparison between the two plots in Figure 3 reveals a significant difference in the representation organization. STMEM exhibits a gradual convergence process, with patch similarity steadily increasing across layers before dropping sharply in the final layers. IPASTMEM achieves a high similarity early on (layers 4–5) and then gradually declines. This suggests that IPASTMEM tends to “compress” information at the beginning of the network, then re-establishes representational diversity in later layers. Such behavior arises in IPASTMEM because all layers contribute equally during pretraining, with the early and middle layers also actively involved in the reconstruction process. Notably, these properties are consistently maintained across all ECG leads, confirming that the model’s representation learning mechanism does not depend on the individual characteristics of each signal channel but instead exhibits a generalization tendency across the entire input data space. Thus, pretrained ViT models in multi-lead ECG signal processing demonstrate high stability and adaptability. The results of the average cosine similarity across layers explain the phenomenon observed in the linear probing evaluation across layers: the middle layers often offer superior performance for downstream tasks. The reason is that the model reaches an optimal information synthesis state at this stage, in which the representations effectively encode the specific spatio-temporal relationship of the ECG signal.
@@ 6.3 Average attention entropy:
Analyzing the average attention entropy (AAE) across layers, as shown in Figure 5, provides additional evidence to explain the performance degradation on downstream classification tasks when using the representation from the last layer. Both models exhibit very high entropy in the first layer, reflecting the dispersion of attention and the high uncertainty about the input information. This is typical of the early stages of representation learning when the model has not yet identified the critical regions in the signal. From the second layer onward, the entropy drops rapidly. It stabilizes at a low level, indicating that the model begins to focus on meaningful information and the representation becomes more certain. In the final layers, the entropy increases slightly again, indicating that the model pays more attention to the fine-grained features in the ECG signal. However, the entropy level at the last layer is still significantly lower than at the first. When these observations are compared with the results from linear probing and cosine similarity, it can be concluded that the final layers are not a crude representation but rather a reconstruction stage designed to focus on fine-grained features. Consequently, performance at the final layers tends to decline compared to the middle layers, yet remains substantially higher than at the early stages. The mathematical definition of AAE is provided in Equation 8, 9.
@@ 7 Experiments and Results
We evaluated the ECG representations learned by the various methods using two popular strategies: linear probing and fine-tuning, applied in both in-distribution and OOD settings as in [34]. Experimental analyses demonstrated that the proposed methods significantly improve most evaluation metrics, indicating superior performance to the baselines.
@@ 7.3 Ablation Study
The combined analysis from Tables VII, VIII, and IX (Appendix E) shows a clear difference in the performance gap between linear probing and full model fine-tuning across the method groups. SimCLR consistently exhibits a huge gap: relatively low performance with linear probing but a sharp increase with fine-tuning on both PTB-XL (in-distribution and OOD) and Chapman, reflecting that the representations learned from SimCLR are not informative enough and rely almost entirely on full parameter updates to achieve high performance. In contrast, CPC and STMEM exhibit only a small gap, indicating stable representation quality and better generalization ability. In particular, the proposed methods maintain a small gap between the two settings and, in some cases, slightly degrade performance, indicating that they have learned informative and stable representations robust enough to achieve high performance with linear probing alone. However, when applying fine-tuning, the proposed methods maintain superior performance over the entire baseline, while the observed characteristics in Tables IV, V, and Table VI in Appendix D (linear probing) remain unchanged.
@@ A.2 Masked Vision Transformer
The ECG patches are randomly masked in the pre-text task. The unmasked patches are fed into the ViT encoder, while the masked patches are reconstructed through the decoder-based Transformers as illustrated in Figure 6.
Encoder: In the pretraining phase, a random masking strategy is applied on the embedding sequence to reduce the inherent redundant information in the ECG signal and simultaneously avoid overfitting. We denote the unmasked embeddings as
@@ Appendix B The proof of Lemma 1
If , then . 25 holds trivially. Consider and then normalizing , :
Applying the norm of both sides of 33 yields
@@ Appendix C Hidden layer analysis
Figure 7 illustrates the classification performance of the STMEM and IPASTMEM models on each representation layer (from layer 1 to layer 12) on four datasets: PTB-XL Conditions, PTB-XL Rhythms, Chapman Conditions, and Chapman Rhythms. The plots show the macro AUC, instance precision, and macro F1 score values when training the classifier on the representations extracted across layers in the pretrained ViT model. The results indicate that the representations at the final layers of the model do not provide the best performance for the downstream classification tasks. The performance improves gradually from the early layers, peaks at the middle layers, and degrades at deeper layers. This trend is consistently observed across multiple evaluation metrics, datasets, and pre-trained models (i.e., STMEM and IPASTMEM), highlighting the generality of the phenomenon. Thus, we can conclude that the middle layers of the Transformer provide richer representations.
@@ Appendix E Ablation study
| Condition | Self-supervised | SimCLR [53, 34] | 84.970.28 92.410.51 (+7.44) | 94.630.31 96.860.18 (+2.23) | 29.371.47 35.341.43 (+5.97) | 97.450.02 97.840.06 (+0.39) | 12.430.62 24.311.23 (+11.88) | 61.880.84 70.280.65 (+8.40) |
| CPC [34] | 89.970.3 91.240.36 (+1.27) | 96.040.09 96.520.09 (+0.48) | 32.600.39 34.610.22 (+2.01) | 97.660.01 97.770.02 (+0.11) | 17.120.69 19.620.69 (+2.50) | 66.230.2 68.260.26 (+2.03) | ||
| STMEM [29] | 92.600.17 93.690.28 (+1.09) | 96.920.07 97.410.1 (+0.49) | 34.780.6 38.261.06 (+3.48) | 97.820.01 98.020.03 (+0.20) | 21.60.9 26.211.43 (+4.61) | 68.410.35 72.40.6 (+3.99) | ||
@@
| Rhythm | Self-supervised | SimCLR [53, 34] | 87.120.28 96.270.6 (+9.15) | 96.280.48 97.320.49 (+1.04) | 81.80.36 89.050.78 (+7.25) | 97.430.04 98.530.08 (+1.1) | 25.821.39 44.882.48 (+19.06) | 83.770.63 89.890.96 (+6.12) |
| CPC [34] | 91.510.37 92.980.74 (+1.47) | 94.330.17 94.580.14 (+0.25) | 82.050.23 83.210.23 (+1.16) | 97.740.03 97.910.04 (+0.17) | 30.692.58 36.52.34 (+5.81) | 82.250.27 83.290.23 (+1.04) | ||
| STMEM [29] | 97.250.21 97.040.26 (-0.21) | 95.330.12 95.510.17 (+0.18) | 85.550.38 86.470.3 (+0.92) | 98.240.05 98.360.04 (+0.12) | 44.12.73 47.872.58 (+3.77) | 85.720.28 86.570.35 (+0.85) | ||
@@
| Condition | Self-supervised | SimCLR [53, 34] | 77.370.35 83.650.44 (+6.28) | 89.440.62 97.70.47 (+8.26) | 36.350.72 54.472.34 (+18.12) | 98.170.02 98.790.08 (+0.62) | 12.520.95 22.591.62 (+10.07) | 57.810.95 79.381.09 (+21.57) |
| CPC [34] | 81.960.48 82.690.34 (+0.73) | 95.190.26 96.130.31 (+0.94) | 47.280.63 49.850.61 (+2.57) | 98.510.02 98.610.03 (+0.1) | 14.870.68 16.560.61 (+1.69) | 70.40.47 73.370.5 (+2.97) | ||
| STMEM [29] | 85.240.22 85.370.25 (+0.13) | 97.760.24 98.220.23 (+0.46) | 53.580.66 55.421.02 (+1.84) | 98.770.01 98.840.03 (+0.07) | 19.761.12 21.820.95 (+2.06) | 78.50.52 80.390.77 (+1.89) | ||
@@
| Rhythm | Self-supervised | SimCLR [53, 34] | 83.310.24 93.870.8 (+10.56) | 86.640.59 97.10.73 (+10.46) | 61.961.24 89.351.28 (+27.39) | 95.550.09 98.510.15 (+2.96) | 37.281.52 55.752.42 (+18.47) | 64.881.07 89.971.43 (+25.09) |
| CPC [34] | 91.470.36 92.390.73 (+0.92) | 94.030.62 95.310.31 (+1.28) | 80.181.25 84.040.75 (+3.86) | 97.320.11 97.780.13 (+0.46) | 45.810.7 47.961.22 (+2.15) | 81.21.12 84.660.8 (+3.46) | ||
| STMEM [29] | 94.420.07 94.40.13 (-0.02) | 97.60.16 97.90.24 (+0.3) | 90.440.47 91.351.02 (+0.91) | 98.670.06 98.750.14 (+0.08) | 57.990.88 59.041.76 (+1.05) | 91.160.29 91.860.87 (+0.7) | ||
@@
| Condition | Self-supervised | SimCLR [53, 34] | 84.650.19 92.380.42 (+7.73) | 94.690.22 96.960.19 (+2.27) | 28.891.13 34.751.51 (+5.86) | 97.470.01 97.850.04 (+0.38) | 13.960.69 24.512.13 (+10.55) | 62.340.31 70.430.73 (+8.09) |
| CPC [34] | 89.70.44 90.640.27 (+0.94) | 95.910.15 96.350.07 (+0.44) | 31.930.25 33.80.31 (+1.87) | 97.620.01 97.740.01 (+0.12) | 15.880.94 19.330.39 (+3.45) | 65.30.39 67.750.22 (+2.45) | ||
| STMEM [29] | 92.170.18 93.560.29 (+1.39) | 96.890.12 97.430.09 (+0.54) | 34.180.36 38.230.75 (+4.05) | 97.790.01 98.020.01 (+0.23) | 21.681.32 27.011.06 (+5.33) | 68.140.31 72.570.43 (+4.43) | ||
@@
| Rhythm | Self-supervised | SimCLR [53, 34] | 84.980.71 96.440.45 (+11.46) | 96.170.36 97.410.43 (+1.24) | 81.380.6 89.090.73 (+7.71) | 97.360.05 98.540.09 (+1.18) | 24.420.87 44.282.87 (+19.86) | 83.320.43 90.060.82 (+6.74) |
| CPC [34] | 90.420.81 92.111.37 (+1.69) | 94.250.11 94.40.07 (+0.15) | 81.280.4 82.220.24 (+0.94) | 97.590.05 97.730.05 (+0.14) | 32.213.04 33.892.37 (+1.68) | 81.540.32 82.370.16 (+0.83) | ||
| STMEM [29] | 96.790.32 96.830.28 (+0.04) | 95.470.11 95.510.2 (+0.04) | 85.780.16 86.490.35 (+0.71) | 98.240.04 98.380.03 (+0.14) | 46.334.18 49.443.26 (+3.11) | 85.980.23 86.690.35 (+0.71) | ||
@@
![[Uncaptioned image]](Figs/PhuNguyen.jpg)
![[Uncaptioned image]](Figs/HuyPhan.jpeg)
![[Uncaptioned image]](Figs/HieuPham.jpg)
![[Uncaptioned image]](x8.png)
![[Uncaptioned image]](Figs/BertVandenberk.jpg)
![[Uncaptioned image]](Figs/MaartenDeVos.png)